A worldwide cosmetics company can upgrade the quality of one of its products by purchasing new equipment at a cost of K155,000. The new equipment would replace old equipment that has a current market value of K23,000. The old equipment originally cost K180,000 and was three quarters depreciated. If the old equipment was used for an additional 12 years, its salvage value at that time would be K5,000. The new equipment has an expected life of 12 years. Its salvage value is estimated at K30,000.
By upgrading the quality of this product, the company would be able to increase the sale price. As a result, the operating income before tax will increase by K20,000 per year for the first 3 years, and by K25,000 per year during the last 9 years.
The company’s tax rate is 40% and its cost of capital after tax is 15%. Depreciation is on straight line basis
Compute the net present value (NPV) and the internal rate of return (IRR) for the investment and state whether the company should proceed with the investment.
1
Expert's answer
2021-09-22T07:50:54-0400

Irr of New equipment=4.06%

Old equipment NPV:

"\frac{-5000}{1.15^{12}}=-934.5"

New equipment NPV:

"-155000+23000+\frac{20000*0.6}{1.15}+\frac{20000*0.6}{1.15^2}+\frac{20000*0.6}{1.15^3}+25000*0.6*\frac{\frac{1}{1.15^4}*(1-\frac{1}{1.15^9})}{1-\frac{1}{1.15}}-\frac{30000}{1.15^{12}}=-63147. 6"

The company should not proceed with the investment